x=y+6
y=-2-x
My daughter is trying to solve this equation by substitution. I am useless in attempting to help her. We would greatly appreciate you explaining the answer.

Answer by checkley77(12844)

x=y+6
NOW SUBSTITUTE (Y+6) FOR X IN THE OTHER EQUATION.
y=-2-x
Y=-2-(Y+6)
Y=-2-Y-6
Y+Y=-8
2Y=-8
Y=-8/2
Y=-4 ANSWER.
X=-4+6
X=2 ANSWER.
PROOF:
-4=-2-2
-4=-4